OpenStudy (anonymous):

Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 Equation #2 8 − 5x = 4x − 1 4x + 4 = 10 − 2x

OpenStudy (anonymous):

8 - 5x = 4x - 1 4x + 4 = 10 - 2x -4x - 5x = -1 - 8 4x - 2x = 10 - 4 -9x = -9 2x = 6 x = 1 x = 3 I need to see your answer choices